1.b given zz = -5 +4i, and zz = 33 + 5i. Find
steps please
Answer #1

Given that = -5+ 4i and zz = 33+5i Then -5+41 33+51 = - 5+41 33 4 5 x 133-59) 43 - ) using (a+b) (9-b) = 9² 6² €5+41)-(33-51)